Given f(x)=x^2+2x+1 find f(x+h) and simplify how can i figure this out if only f is given to me and not h?
Answer by Nate(3500)

f(x) = x^2 + 2x + 1
f(x + h) = (x + h)^2 + 2(x + h) + 1
f(x + h) = x^2 + 2xh + h^2 + 2x + 2h + 1